OpenStudy (anonymous):

1. Does the following equation represent a hyperbola? A. Yes B. No, you cannot take the square root of 3 evenly. C. No, there should be a plus sign in-between the x and y terms. D. No, there are not two squared terms.

OpenStudy (anonymous):

\[\frac{ (x-7)^2 }{ 4 }-\frac{ (y-2) }{ 3 }=1\]

OpenStudy (anonymous):

no, because both of the variables are not squared